# June 13, 2024

### v2

@@ Abstract
This work investigates efficient score-based black-box adversarial attacks with a high Attack Success Rate (ASR) and good generalizability. We design a novel attack method based on a Hierarchical Disentangled Feature space and cross domain, called DifAttack++, which differs significantly from the existing ones operating over the entire feature space. Specifically, DifAttack++ firstly disentangles an image’s latent feature into an adversarial feature (AF) and a visual feature (VF) via an autoencoder equipped with our specially designed Hierarchical Decouple-Fusion (HDF) module, where the AF dominates the adversarial capability of an image, while the VF largely determines its visual appearance. We train such autoencoders for the clean and adversarial image domains respectively, meanwhile realizing feature disentanglement, by using pairs of clean images and their Adversarial Examples (AEs) generated from available surrogate models via white-box attack methods. Eventually, in the black-box attack stage, DifAttack++ iteratively optimizes the AF according to the query feedback from the victim model until a successful AE is generated, while keeping the VF unaltered. Extensive experimental results demonstrate that our method achieves superior ASR and query efficiency than SOTA methods, meanwhile exhibiting much better visual quality of AEs. The code is available at https://github.com/csjunjun/DifAttack.git.
@@ 1 Introduction
It has been observed that meticulously crafted Adversarial Examples (AEs) are beneficial for evaluating and improving the robustness of Deep Neural Networks (DNNs) Dong et al (2020); Croce and Hein (2020b). AEs refer to images that deceive DNNs by incorporating imperceptible perturbations onto clean images. The methods for generating AEs can be roughly classified into three categories based on the information of victim DNNs available to the attacker. The first category is the white-box attack Dong et al (2018); Croce and Hein (2020a); Zhang et al (2021), in which the attack method has access to the architecture and weights of the victim model. The second category is called the black-box attack, where the attacker can only obtain the model’s output. The case where the model’s output contains both categories and scores is known as a score-based black-box attack, while the situation where only categories are included is named a decision-based black-box attack Brendel et al (2018). The third category refers to the gray-box attack Guo et al (2018), where attackers can obtain models’ knowledge that falls between white-box attacks and black-box attacks. Given that contemporary potent classification APIs Google (2024); Imagga (2024) generally provide image prediction categories alongside scores, this work primarily focuses on practical score-based black-box attacks.
@@
We observe that current state-of-the-art (SOTA) score-based black-box attack methods predominantly exhibit two aspects that can be improved. On the one hand, prior techniques that seek AEs by optimizing image features usually operate within the entire feature space, such as AdvFlow Mohaghegh Dolatabadi et al (2020), TREMBA Huang and Zhang (2020) and CGA Feng et al (2022), without embracing the disentanglement of adversarial capability and visual characteristics in the image feature. Optimizing the entire feature of images perhaps results in substantial pixel perturbations that will encounter truncation by the predefined disturbance constraint, thus diminishing the optimization efficacy of the features and eventually increasing query numbers. Another possible chain reaction is the decrease in the Attack Success Rate (ASR) caused by the limited query numbers. Although there are a few black-box methods that achieve leading ASR and query efficiency, the visual quality of their generated AEs is often unsatisfactory and quite perceptible. For instance, SquareAttack Andriushchenko et al (2020) always produce AEs filled with colored stripes and squares. To improve the unsatisfactory ASR and query efficiency, meanwhile maintaining the visual quality of AEs, our DifAttack++ disentangles an image’s latent feature into an Adversarial Feature (AF) and a Visual Feature (VF), where the former dominates the adversarial capability of an image, while the latter largely determines its visual appearance. Afterward, the AF is iteratively optimized according to query feedback, while keeping the VF unaltered.
@@
On the other hand, some score-based attack techniques including P-RGF Cheng et al (2019) and GFCS Lord et al (2022), exploit the gradients from local white-box surrogate models to update AEs during each query. These methods are effective when image categories of the training dataset for surrogate models and the victim model are the same, namely a close-set black-box attack scenario analogy to the close-set recognition setting Vaze et al (2021). When it comes to an open-set black-box attack, wherein the training dataset of the victim model is unknown, the effectiveness of these methods declines, as the possible incongruence between the output categories of surrogate models and the victim model hampers the calculation of gradients from surrogate models. Although there are other approaches, such as SimBA Guo et al (2019a), CGA and SquareAttack, which are workable in both the open-set and close-set scenarios. These methods still face challenges in terms of the query efficiency, ASR or severe degradation of image quality. Our DifAttack++ circumvents the need to compute gradients of surrogate models when updating AEs to be queried. Instead, before embarking upon querying the victim model, we leverage AEs generated by performing white-box attack methods on surrogate models to learn how to extract a disentangled AF and VF from the image’s latent feature, which enables DifAttack++ to more effectively conduct black-box attacks in open-set scenarios. The efficiency of DifAttack++ in open-set scenarios can be attributed to the fact that the disentangled feature space for optimizing AEs is also responsible for the image reconstruction task, which means that an image can still be reconstructed into its original form with minimal loss after being mapped into our disentangled feature space. In contrast, the feature space in other methods, such as TREMBA and CGA, is directly used to generate adversarial perturbations that are tightly bound to the data distribution of the training dataset. The latent feature for image reconstruction in DifAttack++ has better generalizability on unknown datasets than the feature space learned from adversarial perturbation distributions of a specific dataset. In addition, DifAttack++ owns the flexibility to selectively harness surrogate models, e.g., in closed-set scenarios, generating transferable AEs as the starting point for beginning the query attack.
@@
Specifically, in our DifAttack++, we firstly train two distinct autoencoders equipped with our proposed Hierarchical Decouple-Fusion (HDF) module to realize image reconstructions and to disentangle AFs as well as VFs from images’ intermediate features. The feature disentanglement is feasible because the adversarial capability is mainly entwined with the decision boundary of classifiers and is relatively independent of the intrinsic image signal, whereas a contrast holds true for the visual appearance. Here, we use two autoencoders with the same architecture but different weights to extract the features of clean images or their AEs separately, since many existing works Stutz et al (2019); Lin et al (2020) have pointed out that regular AEs are off the manifold of clean images. It has also been verified that the Intrinsic Dimensionality (ID) of AEs are larger than clean images in Qing et al (2024), and data with higher ID are more difficult to be modeled by deep learning Pope et al (2020). Therefore, we disentangle clean images and AEs in cross domain for more precise feature representations.
@@
After obtaining well-trained autoencoders, in the attack stage against victim models, we iteratively optimize AFs of perturbed images using the natural evolution strategy Wierstra et al (2014) according to the query feedback from the victim model, while keeping the VF unaltered, until a successful AE is reconstructed from the fusion of the perturbed AF and the initial image’s VF. The initial image refers to the start point in the attack stage, which is usually set as the clean image in open-set scenarios or a transferable AE in close-set scenarios according to the applicability of surrogate models. More explanations about initial images will be clarified in following Sections. To the best of our knowledge, this is the first work that employs the hierarchical disentangled AFs and VFs in cross domain for black-box adversarial attacks. The contributions of our work can be summarized as follows:
@@
  • We design a new hierarchical disentanglement method in cross domain to distill an AF, which has a significant impact on the image’s adversarial capability while minimizing its influence on visual perception, and a VF that has the opposite attributes, from an image’s latent feature;
@@
  • Based on this disentanglement, we propose a new black-box attack method, DifAttack++, which generates AEs by optimizing the AF according to the query feedback while maintaining the VF invariant.
@@
Note that an earlier version of DifAttack++, called DifAttack Liu et al (2024a), has been published in AAAI2024. Here we conclude the differences of DifAttack++ from this conference version:
  • DifAttack++ designs hierarchical convolution layers for feature disentanglement, considering that one single layer of the decouple-fusion (DF) module in DifAttack may be insufficient to disentangle intensely entangled AFs and VFs.
  • Compared to DifAttack, in which clean images and their AEs share the same antoencoder for image reconstruction and feature disentanglement, the use of cross domain in DifAttack++ achieves a more stable and faster training process, meanwhile leading to better more effective attack performance.
  • DifAttack uses a clean image to start the attack stage in both close-set and open-set scenarios, but DifAttack++ adopts a transferable AE as the start point in close-set scenarios for extracting more powerful AFs, resulting in better attack performance.
The rest of this paper is organized as follows. Section 2 reviews related works on existing score-based black-box attack methods, together with the development of disentangled representation. Section 3 introduces the details of our DifAttack++ attack framework, including the train 3.2 and attack 3.3 two stages. Section 4 shows the performance of DifAttack++ and comparison with the competing attack methods in terms of ASR, query efficiency, and visual quality. Finally, Section 5 concludes.
@@ 2.1 Score-based Black-box Attack Methods
As our work mainly focuses on score-based black-box attack methods, we only conduct literature review on such approaches here.
Score-based black-box attack methods can be broadly categorized into three main classes: transfer-based, query-based, and query-and-transfer-based attacks Feng et al (2022). 1) Transfer-based approaches query the black-box model once using AEs generated on local surrogate models, as explored in Qin et al (2023) and etc. Such approaches take advantage of the transferability of AEs across classifiers with distinct architectures, but their ASRs generally fall behind query-based attack methods. 2) The query-based techniques tackle the black-box optimization through iteratively querying the victim model, including NES Ilyas et al (2018), SignH Al-Dujaili and O’Reilly (2020), $\mathcal{N}$Attack Li et al (2019), SimBA Guo et al (2019a), etc., which exhibit an augmented ASR compared to transfer-based approaches, albeit demanding a greater number of queries. 3) The query-and-transfer-based methods amalgamate the merits of transfer-based and query-based techniques, resulting in an elevated ASR and diminished queries. Certain techniques, such as Subspace Guo et al (2019b), P-RGF Cheng et al (2019), GFCS Lord et al (2022) and BASES Cai et al (2022), update queried AEs by performing white-box attacks against surrogate models with the guidance of the query feedback. Other methods learn adversarial distributions by leveraging pairs of clean images and their AEs obtained by attacking surrogate models, including TREMBA Huang and Zhang (2020) and CGA Feng et al (2022), or approximate the adversarial distribution with the clean data distribution, such as AdvFlow Mohaghegh Dolatabadi et al (2020). Then, the learned feature space for mapping the adversarial distribution is adopted as the search space of AEs. Some other approaches incessantly refine an auxiliary model using the query feedback to directly generate AEs Yin et al (2023a), approximate decision boundaries Yin et al (2023b) or gradients Du et al (2020) of victim models, which are time-consuming and resource-intensive. Notably, MCG Yin et al (2023b) achieves a remarkable ASR and query efficiency by integrating SOTA query-based black-box methods with its proposed meta generator. This generator and surrogate models will be fine-tuned according to query feedback information for adapting to different victim models and test data distributions. More recently, Liu et al (2024b) involves a latest conditional denoising diffusion probabilistic model to learn the transformation from clean images to AEs and then AEs can be directly sampling from this well-trained model.
@@ 2.2 Disentangled Representation
In Bengio et al (2013), it is intuitively defined that Disentangled Representation (DR) separates the distinct, independent and informative generative factors of variation in the data. Recently, a lot of endeavors have embraced the utilization of autoencoders Hinton and Zemel (1993); Lee et al (2020); Li et al (2023), generative adversarial networks Goodfellow et al (2014); Arora et al (2023); Meo et al (2024), generative flows Ma et al (2020), and diffusion-based models Chen et al (2024); Yang et al (2024) to acquire disentangled representations of data features, for which mutual information Chen et al (2016); Higgins et al (2017); Chen et al (2018); Burgess et al (2018); Kim et al (2021); Lin et al (2021) and total correlation Kim and Mnih (2018); Chen et al (2018); Achille and Soatto (2018); Meo et al (2024) are mostly widely used. Regarding the structure of disentangled features, DR methods can be categorized into dimension-wise and vector-wise Wang et al (2022). Dimension-wise DR methods result in each fine-grained generative factor represented by a 1-dimension scalar in the whole feature representation, while vector-wise DR methods lead to several coarse-grained disentangled factors represented by associated vectors, of which the dimension is larger than two. Our DifAttack++ adopts vector-wise features, which are more suitable for addressing complicated practical tasks. In practical, DR learning has been broadly applied to domain adaption Peng et al (2019); Cai et al (2019), image-to-image translation Gonzalez-Garcia et al (2018); Lee et al (2018a, 2020), person re-identification Choi et al (2020); Wei et al (2023), identity swapping Gao et al (2021), zero-shot recognition Tong et al (2019), subject-driven text-to-image generation Chen et al (2024), and etc.
The usage of DR in the adversarial defense and attack. In the adversarial defense and attack domains involved in this work, DR is commonly harnessed for detecting AEs or improving models’ robustness, such as Mustafa et al (2020); Wang et al (2021); Yang et al (2021a, b); Qing et al (2024). Mustafa et al (2020) class-wise disentangles the intermediate feature representations of classifiers in the training stage, specifically forcing the features of each class to lie inside a convex polytope that is maximally separated from the polytopes of other classes. In Wang et al (2021), a one-off combo-variational autoencoder is applied to disentangling latent codes, making one latent code only affect one semantic feature. This autoencoder is then used for detecting and purifying AEs based on feature manipulation. Yang et al (2021a) trains a variational autoencoder with the help of a classifier to decompose an input image into a part retaining only essential information for classification (i.e. class-essential) and a part covering all the rest redundant information (i.e. class-redundant). Then this work detects the input image as an AE if its kernel density Lee et al (2018b) of class-essential output is below a threshold. Yang et al (2021b) focuses on enhancing classifiers’ robustness by training two auxiliary class-specific/-irrelevant classifiers to disentangle class-relevant/-irrelevant features. Qing et al (2024) disentangles the representations of clean counterpart and adversarial perturbations for every input image using several autoencoders, then one of the autoencoders is used for calculate the feature distance between the input image and reconstructed image. Finally, this feature distance is used for detecting whether the input is adversarial.
@@
To the best of our knowledge, only a limited number of white-box attack methods have involved feature disentanglement. RNF Kim et al (2021) is a white-box attack approach that optimizes image perturbations by maximizing the gradient of non-robust features in perturbed images, wherein non-robust features are distilled from the target classifier’s intermediate features based on information bottleneck Alemi et al (2016). Another white-box attack method Lu et al (2021), called SSAE, trains an autoencoder to directly generate target models’ AEs. Its training loss aims to minimize the norm difference of output logits between clean images and AEs to achieve better visual consistency meanwhile maximizing the angle discrepancy of output logits for mis-classification.

### v3

# July 1, 2024

@@ Abstract
This work investigates efficient score-based black-box adversarial attacks with a high Attack Success Rate (ASR) and good generalizability. We design a novel attack method based on a hierarchical DIsentangled Feature space, called DifAttack++, which differs significantly from the existing ones operating over the entire feature space. Specifically, DifAttack++ firstly disentangles an image’s latent feature into an Adversarial Feature (AF) and a Visual Feature (VF) via an autoencoder equipped with our specially designed Hierarchical Decouple-Fusion (HDF) module, where the AF dominates the adversarial capability of an image, while the VF largely determines its visual appearance. We train such two autoencoders for the clean and adversarial image domains (i.e., cross-domain) respectively to achieve image reconstructions and feature disentanglement, by using pairs of clean images and their Adversarial Examples (AEs) generated from available surrogate models via white-box attack methods. Eventually, in the black-box attack stage, DifAttack++ iteratively optimizes the AF according to the query feedback from the victim model until a successful AE is generated, while keeping the VF unaltered. Extensive experimental results demonstrate that our DifAttack++ leads to superior ASR and query efficiency than state-of-the-art methods, meanwhile exhibiting much better visual quality of AEs. The code is available at https://github.com/csjunjun/DifAttack.git.
@@ 1 Introduction
Adversarial Examples (AEs) refer to images that deceive Deep Neural Networks (DNNs) by incorporating imperceptible perturbations onto clean images. Although AEs pose a security threat to many DNNs, it has also been observed that meticulously crafted AEs are beneficial for evaluating and improving the robustness of DNNs Dong et al (2020); Croce and Hein (2020b); Tian et al (2021); Wu et al (2022); Kazemi et al (2023); Li et al (2024), and can even be employed for the privacy protection Liu et al (2023, 2024a). The methods for generating AEs can be roughly classified into three categories based on the information of victim DNNs available to the attacker. The first category is the white-box attack Dong et al (2018); Croce and Hein (2020a); Zhang et al (2021), in which the attacker has access to the architecture and weights of the victim model. The second category is called the black-box attack, where the attacker can only obtain the model’s output. The case where the model’s output contains both categories and scores is known as a score-based black-box attack, while the situation where only categories are included is named a decision-based black-box attack Brendel et al (2018). The third category refers to the gray-box attack Guo et al (2018), where the attacker can obtain models’ knowledge that falls between white-box attacks and black-box attacks. Given that contemporary potent classification APIs, e.g., Google (2024); Imagga (2024), generally provide image prediction categories alongside scores, this work primarily focuses on the practical score-based black-box attacks.
@@
We observe that State-Of-The-Art (SOTA) score-based black-box attack methods predominantly exhibit two aspects that can be improved. On the one hand, prior techniques that seek AEs by optimizing image features usually operate within the entire feature space, such as AdvFlow Mohaghegh Dolatabadi et al (2020), TREMBA Huang and Zhang (2020) and CGA Feng et al (2022), without embracing the disentanglement of adversarial capability and visual characteristics in the image feature. Optimizing the entire feature of images could result in substantial pixel perturbations that will encounter truncation by the predefined distortion constraint, thus diminishing the optimization efficacy of the features and eventually increasing query numbers. Another possible chain reaction is the decrease in the Attack Success Rate (ASR) caused by the limited query numbers. Although there are a few black-box methods that achieve leading ASR and query efficiency, the visual quality of their generated AEs is often unsatisfactory and quite perceptible. For instance, SquareAttack Andriushchenko et al (2020) usually produces AEs filled with colored stripes and squares. To improve the unsatisfactory ASR and query efficiency, meanwhile maintaining the visual quality of AEs, our DifAttack++ disentangles an image’s latent feature into an Adversarial Feature (AF) and a Visual Feature (VF), where the former dominates the adversarial capability of an image, while the latter largely determines its visual appearance. Afterward, the AF is iteratively optimized according to the query feedback, while keeping the VF unaltered.
@@
On the other hand, some score-based attack techniques including P-RGF Cheng et al (2019) and GFCS Lord et al (2022), exploit the gradients from local white-box surrogate models to update AEs during each query. These methods are effective when image categories of the training dataset for surrogate models and the victim model are the same, namely a close-set black-box attack scenario analogy to the close-set recognition setting Vaze et al (2021); Shao et al (2022). When it comes to an open-set black-box attack, wherein the training dataset of the victim model is unknown, the effectiveness of these methods declines dramatically, as the possible incongruence between the output categories of surrogate models and the victim model hampers the calculation of gradients from surrogate models. Although there are other approaches, such as SimBA Guo et al (2019a), CGA Feng et al (2022) and SquareAttack Andriushchenko et al (2020), which are workable in both open-set and close-set scenarios, they still face challenges in terms of the query efficiency, ASR or severe degradation of image quality. In contrast, our DifAttack++ circumvents the need to compute gradients of surrogate models when updating AEs to be queried. Instead, before embarking on querying the victim model, we leverage AEs generated by performing white-box attack methods on surrogate models to learn how to extract a disentangled AF and VF from the image’s latent feature, which enables DifAttack++ to more effectively conduct black-box attacks in open-set scenarios. The efficiency of DifAttack++ in open-set scenarios can be attributed to the fact that the disentangled feature space for optimizing AEs is also responsible for the image reconstruction task, which means that an image can still be reconstructed into its original form with a minimal loss after being mapped into our disentangled feature space. However, the feature space in other methods, such as TREMBA Huang and Zhang (2020) and CGA Feng et al (2022), is directly used to generate adversarial perturbations that are tightly bound to the data distribution of the training dataset. The latent feature for image reconstructions in DifAttack++ has better generalizability to unknown datasets than the feature space learned from adversarial perturbation distributions of a specific dataset. In addition, DifAttack++ owns the flexibility to selectively harness surrogate models, e.g., in closed-set scenarios, generating transferable AEs as the starting point for the query attack.
@@
More specifically, in DifAttack++, we first train two distinct autoencoders equipped with the designed Hierarchical Decouple-Fusion (HDF) module to achieve image reconstructions and to disentangle AFs as well as VFs from images’ latent features. In the training stage, our proposed feature interchange operations in Cross-Domain (CD), i.e., clean and adversarial image domains, also cooperate with HDF in learning disentangled features. The feature disentanglement is feasible because the adversarial capability is mainly entwined with the decision boundary of classifiers and is relatively independent of the intrinsic image signal, whereas a contrast holds true for the visual appearance. Here, we use two autoencoders with the same architecture but different weights to extract the features of clean images and their AEs separately, since many existing works Stutz et al (2019); Lin et al (2020) have pointed out that regular AEs are off the manifold of clean images. It has also been verified that the Intrinsic Dimensionality (ID) of AEs is larger than that of clean images in Qing et al (2024), and data with higher ID are more difficult to be modeled by DNNs Pope et al (2020). Therefore, we reconstruct and disentangle clean images and AEs in CD for more precise feature representations.
@@
After obtaining these well-trained autoencoders, in the attack stage against victim models, we iteratively optimize AFs of perturbed images using the natural evolution strategy Wierstra et al (2014), according to the query feedback from the victim model, while keeping the VF unaltered, until a successful AE is reconstructed from the fusion of the perturbed AF and the initial image’s VF. Here, the initial image refers to the starting point in the attack stage, which is usually set as the clean image in open-set scenarios or a transferable AE in close-set scenarios, depending on the applicability of surrogate models. More explanations about initial images will be clarified in the following sections. To the best of our knowledge, this is the first work that employs the hierarchical disentangled AFs and VFs in CD for black-box adversarial attacks. The contributions of our work can be summarized as follows:
@@
  • We design a new hierarchical disentanglement method to distill an AF, which has a significant impact on the image’s adversarial capability while minimizing its influence on visual perception, and a VF that has the opposite attributes, from an image’s latent feature;
@@
  • Based on this disentanglement, we propose a new query-efficient score-based black-box attack method, DifAttack++, which generates AEs by optimizing the AF according to the query feedback while maintaining the VF invariant;
@@
Note that an earlier version of DifAttack++, called DifAttack, has been published in AAAI2024 Liu et al (2024b). Below, we summarize the differences of DifAttack++ from its conference version:
  • DifAttack++ designs hierarchical convolution layers for the feature disentanglement, considering that one single layer of the Decouple-Fusion (DF) module in DifAttack is insufficient to disentangle intensely entangled AFs and VFs;
  • Compared to DifAttack, in which clean images and their AEs share the same autoencoder for the image reconstruction and feature disentanglement, the use of two cross-domain autoencoders in DifAttack++ achieves a more stable and faster training process, meanwhile leading to more effective black-box attack performance;
  • DifAttack uses a clean image to start the attack stage in both close-set and open-set scenarios; but DifAttack++ adopts a transferable AE as the starting point in close-set scenarios for extracting more powerful AFs, resulting in better attack efficiency;
  • The ASR and query efficiency of DifAttack++ far surpass those of DifAttack, e.g., average 17.2% higher ASR and 53.7% lower average query numbers, when performing targeted attacks on ImageNet against four complex victim models in closed-set scenarios.
The rest of this paper is organized as follows. Section 2 reviews related works on existing score-based black-box attack methods, together with the development of disentangled representation. Section 3 introduces the details of our DifAttack++ attack framework, including the training (Section 3.1) and attack (Section 3.2) stages. Section 4 shows the performance of DifAttack++ and comparisons with competing attack methods in terms of ASR, query efficiency, and visual quality. Finally, Section 5 gives a conclusion.
@@ 2.1 Score-based Black-box Attack Methods
As our work mainly focuses on score-based black-box attack methods, we now conduct a literature review on such approaches.
Score-based black-box attack methods can be broadly categorized into three main classes: transfer-based, query-based, and query-and-transfer-based attacks Feng et al (2022). Transfer-based approaches query the black-box model once using AEs generated on local surrogate models, as explored in Qin et al (2023) and etc. These approaches take advantage of the transferability of AEs across classifiers with distinct architectures; but their ASRs generally are not satisfactory. The query-based techniques tackle the black-box optimization through iteratively querying the victim model. Some representatives are NES Ilyas et al (2018), $\mathcal{N}$Attack Li et al (2019), SimBA Guo et al (2019a), SignH Al-Dujaili and O’Reilly (2020), etc., which exhibit improved ASR performance compared to transfer-based approaches, albeit demanding a greater number of queries. Furthermore, the query-and-transfer-based methods amalgamate the merits of transfer-based and query-based techniques, resulting in an elevated ASR and smaller numbers of queries. For instance, Subspace Guo et al (2019b), P-RGF Cheng et al (2019), GFCS Lord et al (2022) and BASES Cai et al (2022) update queried AEs by performing white-box attacks against surrogate models with the guidance of the query feedback. There are some other methods that learn adversarial distributions by leveraging pairs of clean images and their AEs obtained by attacking surrogate models, e.g., TREMBA Huang and Zhang (2020) and CGA Feng et al (2022), or approximate the adversarial distribution with the clean data distribution, such as AdvFlow Mohaghegh Dolatabadi et al (2020). Then, the learned feature space for mapping the adversarial distribution is adopted as the search space of AEs. We also observe that some approaches incessantly refine auxiliary models using the query feedback to directly generate AEs Ma et al (2021), approximate decision boundaries Yin et al (2023) or gradients Du et al (2020) of victim models. However, these methods are often time-consuming and resource-intensive. Notably, MCG Yin et al (2023) achieves a remarkable ASR and query efficiency by integrating SOTA query-based black-box methods with their proposed meta generator. This generator and surrogate models will be fine-tuned according to the query feedback for adapting to different victim models and test data distributions.
@@ 2.2 Disentangled Representation
As this work primarily focuses on attacking in the disentangled feature space, we also would like to briefly introduce the development of Disentangled Representation (DR), followed by a short review of adversarial attacks incorporating DR.
@@
In Bengio et al (2013), it is intuitively defined that DR separates the distinct, independent and informative generative factors of variation in the data. Recently, many endeavors have embraced the utilization of autoencoders Lee et al (2020); Li et al (2023), generative adversarial networks Meo et al (2024), generative flows Ma et al (2020), and diffusion-based models Chen et al (2024); Yang et al (2024) to acquire disentangled representations of data features, for which mutual information Burgess et al (2018); Kim et al (2021); Lin et al (2021) and total correlation Kim and Mnih (2018); Achille and Soatto (2018); Meo et al (2024) are mostly widely used. Regarding a more comprehensive discussion for DR, we refer interested readers to Wang et al (2022). DR has been broadly applied to the research on domain adaption Peng et al (2019); Cai et al (2019), image-to-image translation Gonzalez-Garcia et al (2018); Lee et al (2020), person re-identification Wei et al (2023), and pose estimation Nie and Liu (2021) etc.
@@
As for the areas on adversarial attack and its defense strategies, DR has been commonly harnessed for detecting AEs or improving models’ robustness; see Mustafa et al (2020); Wang et al (2021); Yang et al (2021); Qing et al (2024). To the best of our knowledge, only a limited number of adversarial attack methods have involved DR, and all of them are white-box approaches. Along this line, RNF Kim et al (2021) optimizes image perturbations by maximizing the gradient of non-robust features in perturbed images, wherein non-robust features are distilled from the target classifier’s intermediate features. Another white-box attack method Lu et al (2021), called SSAE, trains an autoencoder to directly generate target models’ AEs. SSAE aims to minimize the norm difference of output logits between clean images and AEs to achieve better visual consistency, meanwhile maximizing the angle discrepancy of output logits for mis-classification.
Note that the disentanglement strategy of our DifAttack++ is quite different from the existing ones. White-box attack methods generally have prior knowledge of the victim model’s parameters and architecture during the DR learning, while DifAttack++ relies on surrogate models to learn DRs, with the particular challenging requirement of the generalizability of DRs across diverse classifiers. Furthermore, DifAttack++ disentangles image features capable of reconstructing the image signals, rather than the ones extracted by specific classifiers as Mustafa et al (2020); Kim et al (2021) do. Also, none of the existing methods disentangles adversarial capability and visual perception from latent features. As will be clear soon, the features disentangled by DifAttack++ are not closely related to specific categories or semantics, making DifAttack++ exhibit much more powerful disentanglement capability in open-set scenarios.
3 Proposed DifAttack++
Our DifAttack++ aims to use disentangled VFs and AFs of images to decode AEs in score-based black-box settings, guided by the query feedback from the victim model. As illustrated in Fig. 1, DifAttack++ unfolds in two phases. In the training stage, we train two autoencoders $\mathcal{G}$ and $\mathcal{G}^{*}$ with the same model architecture but different weights. These two autoencoders both are able to conduct feature disentanglement and image reconstruction. The training images of $\mathcal{G}$ and $\mathcal{G}^{*}$ are a set of clean image $\mathbf{x}$ and its AE $\mathbf{x}^{*}$, where $\mathbf{x}^{*}$ is generated by performing an untargeted white-box attack on surrogate models. Taking $\mathbf{x}$ as an example, $\mathcal{G}$ can disentangle the latent feature $\mathbf{z}$ of $\mathbf{x}$ into the AF $\mathbf{z}_{a}$ and VF $\mathbf{z}_{v}$ via our proposed HDF module, and then reconstructs another image $\tilde{\mathbf{x}}$ similar to $\mathbf{x}$ by decoding the fusion of $\mathbf{z}_{a}$ and $\mathbf{z}_{v}$. Likewise, $\mathcal{G}^{*}$ also has the same functionality for its input.
In the attack stage, we leverage the pre-trained $\mathcal{G}$ and $\mathcal{G}^{*}$ to generate an AE $\mathbf{x}^{\prime}$ of $\mathbf{x}$ against the black-box victim model $\mathcal{F}$. Specifically, $\mathbf{x}^{\prime}$ is obtained through an iterative optimization process, wherein we continuously sample a batch of perturbed images to extract their AFs $\hat{\mathbf{z}}^{i}_{a}$, while keeping the VF $\hat{\mathbf{z}}_{v}$, which is used for the subsequent image reconstruction along with $\hat{\mathbf{z}}^{i}_{a}$, consistent with that of the initial image. Such a process produces batches of reconstructed images $\hat{\mathbf{x}}^{i}_{f}$ to query $\mathcal{F}$, until either the query number reaches its maximum threshold $Q$, or one of the perturbed images $\hat{\mathbf{x}}^{i}_{f}$ successfully deceives $\mathcal{F}$, namely, $\mathbf{x}^{\prime}=\hat{\mathbf{x}}_{f}^{i}$.
For clarity, a summary of notations frequently used in the following sections is provided in Table 1.
3.1 Training Stage: Train Autoencoders $\mathcal{G}$ and $\mathcal{G}^{*}$
Let $\mathbf{x}$ be a clean image in the training dataset $\mathbb{T}$ and $\mathbb{C}$ be the model set formed by the pre-trained white-box surrogate models based on $\mathbb{T}$. Note that these white-box surrogate models can be readily obtained and their availability has also been assumed in the existing approaches Lord et al (2022); Feng et al (2022). We adopt the white-box attack method PGD Madry et al (2018) to generate an AE $\mathbf{x^{*}}$ of $\mathbf{x}$, by attacking a surrogate model $\mathcal{C}_{j}$. This $\mathcal{C}_{j}$ is selected randomly from $\mathbb{C}$. Clearly, PGD can be replaced by other white-box attack methods or a collection of them (see the supplementary file of DifAttack Liu et al (2024b) for more details). Different from our previous work DifAttack which shares one autoencoder for both clean and adversarial images, in this work, we assign distinct autoencoders $\mathcal{G}$ and $\mathcal{G}^{*}$ for the clean and adversarial domains. Noticing that regular AEs usually lie off the manifold of normal images Lin et al (2020), the assignment of two autoencoders is more reasonable, and could lead to better training efficiency and black-box attack performance, as will be verified experimentally. Below, let us give the details on the disentanglement and reconstruction tasks.